Tell whether the pH at equivalence point is less than 7, greater than 7 or equal to 7 for each of the following solutions. PLEASE SHOW WORK IF POSSIBLE OR EXPLAINATION; Thank you so much!

12. 0.1M acetic acid with 90.10M NaOH

a). Greater than 7

b.) Equal to 7

c.) Less than 7

d.) Equal to 14

e.) Don't Know

13. 0.050M formic acid (HCOOH) with 0.0930M NaOH

a.) Greater than 7

b.) Less than 7

c.) Equal to 7

d.) Equal to 1.0

e.) Equal to 9

14. 0.050M NaOH with 0.50M HCl

a). Equal to 7

b.) Greater than 7

c.) Less than 7

d.) Equal to 14

e.) Equal to 1.0

15. 0.2M HCl with 0.2M NH3

a.) Less than 7

b.) Equal to 7

c.) Greater than 7

d.) Equal to 7

e.) Equal to 0.0
Answer #1

12. CH3COOH + NaOH ----------> CH3COONa + H2O

I       0.1               90.1                         0

C       -0.1           -0.1                          0.1

E      0               90

[OH-]    = 90M

POH = -log[OH-]

          = -log90   = -1.9542

PH   = 14-POH

      = 14-(-1.9542)   = 15.9542>>>>equalt ot 14

d.) Equal to 14

13. HCOOH + NaOH -------------> HCOONa + H2O

I        0.05       0.093                         0

C       -0.05    -0.05                           0.05

E        0          0.043 0.05

[OH-] = 0.043M

POH = -log[OH-]

   = -log0.043     = 1.3665

PH   = 14-POH

         = 14-1.3665   = 12.6335

a.) Greater than 7

14 acid conc is more

c.) Less than 7

15. NH4+ +H2O ---------> NH4OH + H+

I      0.2                                  0              0

C     -x                                   +x             +x

E     0.2-x                              +x               +x

    Ka   = {NH4OH][H+]/[NH4+]

   5.6*10^-10   = x*x/0.2-x

5.6*10^-10*(0.2-x) = x^2

     x = 1.06*10^-5

[H+] = x= 1.06*10^-5

PH = -log[H+]

         = -log1.06*10^-5   = 4.9746

a.) Less than 7